# MILD: Multi-Layer Diffusion Strategy for Complex and Precise Multi-IP Aware Human Erasing

### v1

@@ 3.1 MILD Backbone: Layered LoRA Diffusion
where $\lambda_{t}$ gradually increases to stabilize training. We adopt a staged strategy, first optimizing the background branch before introducing foreground supervision through ramp-up.
During inference, MILD produces composable outputs:
$\bigl{\{}\text{Layer}_{1},\ldots,\text{Layer}_{N},\text{Background}\bigr{\}},$ (5)
where each $\text{Layer}_{k}$ represents an individual instance $x_{k}$. These layers can be selectively combined to reconstruct arbitrary subsets by removing the instance in $\mathcal{R}\subseteq\{1,2,\ldots,N\}$ as:
where $\mathcal{R}$ denotes the set of instances to remove. For each $k\in\mathcal{R}$, the corresponding layer is excluded from the composite and automatically filled by the background layer $x_{\text{bg}}$. This layered formulation supports controllable recomposition and improves reconstruction fidelity through joint foreground-background supervision.
@@
14:   $\mathcal{L}_{k}\leftarrow\|(1-M_{k})\odot(\epsilon-\epsilon_{k})\|_{2}^{2}$ {Foreground branch}
@@
3.2 Human Morphology Guidance
Although the dedicated Layered LoRA diffusion backbone lays the foundation for disentangling instance layers, fine-grained instance awareness remains essential for accurately removing targets and preserving body structures in spatially overlapping scenarios. To handle such complex scenarios, we introduce Human Morphology Guidance (HMG), a plug-and-play module that introduces fine-grained, human-centric priors into the denoising process.
As illustrated in Algorithm 1, HMG is seamlessly integrated into the MILD backbone by embedding pose and parsing information into the conditioning signals for both foreground and background branches. Specifically, preprocessed posure and parsing maps are separately encoded, flattened, linearly projected, and concatenated to form comprehensive human features, as shown in Figure 2 (b).
To further enhance spatial reasoning, we introduce a lightweight mask-based conditioning scheme that explicitly models neighboring instances. For each instance $k$, a context mask is constructed as:
$M^{(k)}=\text{clamp}\Bigl{(}\sum_{j\neq k}M_{j},\,0,1\Bigr{)},$ (7)
which captures the spatial layout of surrounding instances, providing instance-aware guidance for precise IP localization. This mask is encoded through a shallow convolutional layer and concatenated with the target’s mask embedding and human features. Together, these spatially-aware priors provide stronger guidance for accurately separating instances, particularly in occluded scenarios.
By integrating explicit morphological and spatial cues, HMG significantly enhances instance-level disambiguation, especially in challenging scenes involving occlusions, uncommon poses, or partially visible subjects.
Figure 3: Illustration of the proposed Spatially-Modulated Attention (SMA) mechanism. For each query-key pair $(i,j)$, an spatial bias $\alpha_{st}$ is applied to the vanilla attention score $A_{ij}$, where $s=m_{i}$ and $t=m_{j}$ indicate their foreground/background status.
3.3 Spatially-Modulated Attention
The original attention mechanisms utilize the entire context to generate reasonable background, but this leads to semantic leakage due to the unrestricted interactions between foreground and background tokens. To mitigate this, we propose Spatially-Modulated Attention (SMA), a lightweight module that enforces spatially constrained attention based on mask boundaries, as shown in Figure 3.
In standard self-attention, logits are computed as:
$A_{ij}=\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}.$ (8)
To control attention flow across semantic separated regions, SMA introduces a learnable bias based on the binary masks $m_{i}$ and $m_{j}$ of tokens $i$ and $j$:
where $\alpha_{st}$ are trainable parameters that modulate attention flow between foreground and background regions. The indicator function $\mathds{1}[m_{i}=s,\,m_{j}=t]$ evaluates to $1$ when token $i$ belongs to region $s$ and token $j$ to region $t$, and 0 otherwise. For instance, $\alpha_{10}$ controls how effective a foreground token attends to the background.
Initialized with zero, these biases preserve the vanilla attention at the start and enabling adaptive spatial modulation during training. This enables explicit control over cross-region interactions without altering the attention mechanism itself. With this selective suppression, the attention flow learns to “attend to what it should” and mitigates leakage across masked boundaries. Integrated with MILD backbone and HMG, SMA improves the semantic isolation and artifacts removal in challenging multi-IP scenes.
Figure 4: Qualitative results produced by MILD (ours) and other methods in real-world scenes. The masked regions (in red) and the corresponding removal results (in green) are highlighted for comparison.
4 Experiments
Method MILD Dataset OpenImages Dataset
[…]
Inst-Inpaint 11.42 0.4100 0.7400 0.8207 23.75
LaMa 35.02 0.164 0.8909 0.8872 24.13 10.38 0.0470 0.9192 0.9293 31.83
Table 1: Quantitative comparison of MILD and other methods on MILD and OpenImages datasets. Best results are bold, second best are underlined. Inst-Inpaint is excluded from the evaluation on MILD dataset as it lacks mask guidance, making it unsuitable for precise localization in multi-IP scenarios.
4.1 Experimental Setup
Datasets
We conduct comprehensive evaluations on two datasets to assess both task-specific performance and cross-domain generalization: 1) Our MILD Dataset: We construct a high-quality dataset specifically tailored for multi-IP erasure. The dataset comprises $10,000$ image pairs featuring complex human interactions, diverse body poses, challenging occlusions, and a wide range of background contexts. The detailed information is introduced in Appendix A. 2) OpenImages Dataset: To evaluate generalization beyond human-centric scenarios, we randomly sample 1,000 images from OpenImages V5 Kuznetsova et al. (2020), covering diverse object categories and photographic conditions.
Implementation Details
We train MILD LoRA on 20,00020{,}000 images, including 10,00010{,}000 real-world pairs from our curated MILD dataset and 10,00010{,}000 synthetic samples generated by compositing segmented humans onto OpenImages backgrounds with random cropping and mask augmentation. The baseline model Podell et al. (2023) is optimized using AdamW Loshchilov and Hutter (2019) with a learning rate of 1e1e66, batch size $4$, and trained for 20,00020{,}000 iterations. Human Morphology Guidance leverages pose maps from OpenPose Cao et al. (2017) and parsing maps from SCHP Li et al. (2020b), preprocessed prior to training. A staged training scheme is adopted: the background branch is trained first, followed by a linear ramp-up of foreground losses after 8,0008{,}000 steps. The model employs a UNet backbone within a latent diffusion framework, equipped with rank-$16$ LoRA modules (scaling factor $16$). Inference uses $25$ DDIM steps at $512\times 512$ resolution. All experiments are conducted on NVIDIA $A800$ GPUs.
Baseline Comparisons
We compare against a diverse set of state-of-the-art object removal methods, including SDXL Inpainting Podell et al. (2023), LaMa Suvorov et al. (2022), PowerPaint Zhuang et al. (2024), Inst-Inpaint Yildirim et al. (2023), CLIPAway Ekin et al. (2024), and RoRem Li et al. (2025), covering paradigms from mask-guided inpainting to instruction-based and vision-language-guided editing. To ensure a fair comparison, we adopt the official open-source implementations and reproduce results under a unified evaluation pipeline. All models are either used as publicly released or retrained with default settings on our datasets to ensure fair comparison.
Evaluation Metrics
We adopt five complementary metrics to evaluate object removal quality from perceptual, semantic, and low-level fidelity perspectives. FID Heusel et al. (2017) evaluates global visual realism by comparing feature distributions of generated and real images, while LPIPS Zhang et al. (2018) focuses on local perceptual similarity based on deep features. DINO cosine similarity Caron et al. (2021) and CLIP Image Similarity Radford et al. (2021) measure semantic consistency by computing cosine similarity between the embeddings of generated and ground-truth images, capturing structural and high-level semantic alignment, respectively. Finally, PSNR Hore and Ziou (2010) quantifies pixel-level fidelity within the removed regions.
4.2 Quantitative and Qualitative Comparisons
Table 1 summarizes the quantitative performance of various object removal methods on the MILD and OpenImages datasets. Our proposed MILD strategy consistently outperforms all baselines on the more challenging MILD dataset and remains highly competitive on OpenImages. Compared to the SDXL Inpainting baseline, MILD reduces FID and LPIPS by $54.53\%$ and $61.57\%$ respectively, reflecting significant improvement in visual realism and perceptual quality. It also achieves higher DINO, CLIP, and PSNR scores, demonstrating enhanced semantic consistency and pixel-level fidelity through its spatially-aware layered diffusion design. While LaMa performs well on FID and LPIPS in relatively simple scenes due to its smooth outputs, it underperforms in semantic metrics like DINO and CLIP, revealing limitations in understanding complex content. In contrast, MILD delivers superior results across all dimensions, particularly in complex human interactive scenes.
While traditional quantitative metrics provide valuable evaluations of model performance, they often fail to capture the direct naturalness and visual quality as perceived by human observers. To address this limitation, we conduct a comprehensive perceptual evaluation from both AI and human perspectives. The evaluation of both assessments are summarized in Figure 6, and the detailed experiment setting and the full table is shown in Appendix. Both AI and human evaluations consistently demonstrate that MILD achieves superior perceptual quality compared to baseline methods. Our method (MILD) achieves the highest success rate for both AI $(82\%)$ and human raters $(78\%)$, indicating consistent perceptual superiority. This comprehensive perceptual assessment validates the effectiveness of our approach in producing human-preferred removal results.
Figure 5: Qualitative ablation comparison of our method.
4.3 Ablation Study
The SDXL Inpainting baseline performs poorly on the MILD dataset, often produces incomplete removals or hallucinated content when handling large or occluded subjects. Introducing MILD Backbone significantly improves both perceptual and semantic quality by disentangling generation into background and instance-specific foreground branches. This reduces the repainting phenomenon with 45.7145.71% improvement in FID metric. Building on this backbone, HMG and SMA provide complementary enhancements. As shown in Figure 5, HMG helps to preserve body structure (row $2$) and relieve the residual artifacts along object boundaries(row $1$). Meanwhile, SMA focuses on refining the attention flow, effectively suppressing the unnatural mask bleeding and background leakage, resulting in smoother transitions and more coherent backgrounds. The full MILD model combines these advantages, achieving the best overall performance across all metrics. These results validate the non-redundant and effectiveness of our modular design.
[…]
SDXL Inpaint 53.45 0.242 0.8398 0.8331 18.22
+ MILD Backbone 29.02 0.142 0.9110 0.9025 23.45
   + HMG 26.10 0.120 0.9440 0.9312 24.67
   + SMA 25.10 0.105 0.9380 0.9227 25.21
MILD (Full) 24.20 0.093 0.9703 0.9499 26.24
Table 2: Ablation study on the MILD dataset.
4.4 Compositional Flexibility of Layered Outputs
Beyond achieving high-quality foreground removal and background restoration, MILD’s layered diffusion design introduces a key advantage: the ability to flexibly recompose scenes using instance-specific foreground layers and a clean background. As shown in Figure 7, MILD generates disentangled outputs for each target, enabling selective combination to construct customized scenes. This allows users to separate specific individuals (e.g., only IP11), remove arbitrary targets (e.g., w/o IP33), or reconstruct any subset of the original scene. Such instance-aware recomposition offers fine-grained control over content manipulation, making MILD well-suited for interactive editing and scene simulation tasks. More details are shown in Appendix D.
Figure 7: Illustration of MILD’s compositional flexibility.
5 Conclusion
We present Multi-Instance Layered Diffusion, a diffusion-based strategy reformulating inpainting as a layered, decoupled generation process for precise human erasing, supported by our curated, high-quality MILD dataset. MILD leverages HMG to provide instance-aware structure understanding, and employs SMA to bring spatially-conditioned attention control, effectively reducing artifacts and improving reconstruction fidelity. Extensive experiments show that MILD outperforms existing inpainting baselines, achieving better semantic consistency, fewer artifacts, and more coherent scene reconstructions. MILD provides a promising direction for scaling precise diffusion-based inpainting to complex multi-object and open-domain removal tasks.
Appendix
Appendix A Dataset and Code Availability
Dataset Construction.
Our dataset comprises high-quality images collected from two primary sources: (1) a carefully curated subset of the OpenImages V5 dataset Kuznetsova et al. (2020), and (2) publicly available web images, gathered under fair use for academic research purposes. All images are manually verified to ensure diversity in scenes, human poses, occlusions, and background complexity.
Each image is annotated with pixel-level, instance-specific human masks. We further provide human-verified ground-truth completions to support supervised evaluation of object removal performance. Representative samples are shown in Figure 8. The complete dataset will be publicly released upon acceptance at: https://mild-multi-layer-diffusion.github.io/dataset/
Figure 8: Sample data collected from MILD dataset.
Code and Model Release.
To support future research and ensure reproducibility, we have created a public project page featuring representative examples and the paper abstract. The full codebase and pre-trained models will be made available upon acceptance at: https://mild-multi-layer-diffusion.github.io/
Appendix B More Related Works
Layered Generation and Controllability.
Recent works on image generation have explored layered modeling and layout-based generation to enhance object awareness and controllability. Layered modeling works Jia et al. (2024); Lu et al. (2020) leverage multi-level structures through latent decomposition and scene-layer separation to enable semantically consistent editing. However, they often operate within a unified generation pathway and lack explicit instance disentanglement or spatial conditioning, making them less suitable for multi-object removal. Prompt- and layout-controllable generation works Liu et al. (2022); Zheng et al. (2023); Zhangli et al. (2024) aim to synthesize scenes from object-level prompts or layout priors. Although these methods offer flexible conditioning and structural guidance, they generally operate at a holistic level without explicit separation of foreground instances, limiting their capacity to handle occlusions or support fine-grained instance control. In contrast, our approach introduces a mask-aware, multi-branch generation framework that explicitly disentangles each foreground instance and the background into independent pathways. This layered formulation enables precise object removal, instance-level recomposition, and enhanced semantic fidelity through joint supervision.
Appendix C More Framework Details
To ensure reproducibility, we provide additional details of the training pipeline, architectural design, and spatial supervision strategies used in LayeredUNet. The model is optimized in stages (Algorithm 2), where a training controller gradually activates different generation branches. Each image is encoded into the latent space and perturbed with spatial offsets derived from the union of all masks. To stabilize learning, the background branch is initially protected by freezing its gradients, while foreground branches are progressively optimized with instance-specific conditioning.

### v2

@@ 3.5 Optimization Objective
where $s$ is the training iteration index, and $\theta_{\mathrm{fg}}$ and $\theta_{\mathrm{bg}}$ refer to the foreground and background parameters respectively. We adopt a staged strategy, first optimizing the background branch before ramping in foreground supervision.
Figure 5: Qualitative results produced by MILD (ours) and other methods in real-world scenes. The masked regions (in red) and the corresponding removal results (in green) are highlighted.
4 Experiments
Method MILD Dataset OpenImages Dataset
[…]
Inst-Inpaint 11.42 0.4100 0.7400 0.8207 23.75
LaMa 35.02 0.164 0.8909 0.8872 24.13 10.38 0.0470 0.9192 0.9293 31.83
Table 1: Quantitative comparison of MILD and other methods on MILD and OpenImages datasets. Best results are bold, second best are underlined. Inst-Inpaint is excluded from evaluation on MILD dataset as it lacks mask guidance, making it unsuitable for precise localization in multi-IP scenarios.
4.1 Experimental Setup
Datasets
We conduct comprehensive evaluations on two datasets to assess both task-specific performance and cross-domain generalization: 1) Our MILD Dataset: We construct a high-quality dataset specifically tailored for multi-IP erasure. The dataset comprises $10,000$ image pairs featuring complex human interactions, diverse body poses and challenging occlusions. The detailed information is introduced in Appendix A. 2) OpenImages Dataset: To evaluate generalization beyond human-centric scenarios, we randomly sample 1,000 images from OpenImages V5 (openimages2020), covering diverse object categories and photographic conditions.
Baseline Comparisons
We compare against a diverse set of state-of-the-art object removal methods, including SDXL Inpainting (podell2023sdxl), LaMa (suvorov2022lama), PowerPaint (powerpaint2024), Inst-Inpaint (instinpaint2023), CLIPAway (clipaway2023), and RoRem (rorem2025), covering paradigms from mask-guided inpainting to instruction-based and vision-language-guided editing. To ensure a fair comparison, we adopt the official open-source implementations and reproduce results under a unified evaluation pipeline. All models are either used as publicly released or retrained with default settings on our datasets to ensure fair comparison.
Evaluation Metrics
We adopt five complementary metrics to evaluate object removal quality from perceptual, semantic, and low-level fidelity perspectives. FID (heusel2017gans) evaluates global visual realism by comparing feature distributions of generated and real images, while LPIPS (zhang2018unreasonable) focuses on local perceptual similarity based on deep features. DINO cosine similarity (caron2021emerging) and CLIP Image Similarity (radford2021learning) measure semantic consistency by computing cosine similarity between the embeddings of generated and ground-truth images, capturing structural and high-level semantic alignment, respectively. Finally, PSNR (hore2010image) quantifies pixel-level fidelity within the removed regions.
4.2 Quantitative and Qualitative Comparisons
Table 1 summarizes the quantitative performance of various object removal methods on the MILD and OpenImages datasets. Our proposed MILD strategy consistently outperforms all baselines on the more challenging MILD dataset and remains highly competitive on OpenImages. Compared to the SDXL Inpainting baseline, MILD reduces FID and LPIPS by $54.53\%$ and $61.57\%$ respectively, reflecting significant improvement in visual realism and perceptual quality. It also achieves higher DINO, CLIP, and PSNR scores, demonstrating enhanced semantic consistency and pixel-level fidelity through its spatially-aware layered diffusion design.
While LaMa performs well on FID and LPIPS in relatively simple scenes due to its smooth outputs, it underperforms in semantic metrics like DINO and CLIP, revealing limitations in understanding complex content. In contrast, MILD delivers superior results across all dimensions, particularly in complex human interactive scenes.
While traditional quantitative metrics provide valuable evaluations of model performance, they often fail to capture the direct naturalness and visual quality as perceived by human observers. To address this limitation, we conduct a comprehensive perceptual evaluation from both AI and human perspectives. The evaluation of both assessments are summarized in Figure 6, and the detailed experiment setting and the full table is shown in Appendix G.5. Both AI and human evaluations consistently demonstrate that MILD achieves superior perceptual quality compared to baseline methods. Our method (MILD) achieves the highest success rate for both AI $(82\%)$ and human raters $(78\%)$, indicating consistent perceptual superiority. This comprehensive perceptual assessment validates the effectiveness of our approach in producing human-preferred removal results.
Figure 7: Qualitative ablation comparison of our method.
4.3 Ablation Study
The single layer baseline performs poorly(row $1$) on the MILD dataset, often produces incomplete removals or hallucinated content when handling large or occluded subjects. Introducing MILD Backbone significantly improves both perceptual and semantic quality by disentangling generation into background and instance-specific foreground branches. This reduces the repainting phenomenon with 45.7145.71% improvement in FID metric. Building on this backbone, HMG and SMA provide complementary enhancements. As shown in Figure 7, HMG helps to preserve body structure (row $2$) and relieve the residual artifacts along object boundaries(row $1$). Meanwhile, SMA focuses on refining the attention flow, effectively suppressing the unnatural mask bleeding and background leakage, resulting in smoother transitions and more coherent backgrounds. The full MILD model combines these advantages, achieving the best overall performance across all metrics. These results validate the non-redundant and effectiveness of our modular design.
[…]
53.45 +29.25 0.242 +0.149 0.8398 -0.1305 0.8331 -0.1168 18.22 -8.02
29.02 +4.82 0.142 +0.049 0.9110 -0.0593 0.9025 -0.0474 23.45 -2.79
26.10 +1.90 0.120 +0.027 0.9440 -0.0263 0.9312 -0.0187 24.67 -1.57
25.10 +0.90 0.105 +0.012 0.9380 -0.0323 0.9227 -0.0272 25.21 -1.03
24.20 0.093 0.9703 0.9499 26.24
Table 2: Ablation study on the MILD dataset comparing the SDXL baseline, the model with the proposed MILD backbone, and the incremental addition of HMG and SMA.
5 Discussion
Figure 8: Illustration of MILD’s compositional flexibility.
5.1 Why Does MILD Suppress Frequency Peaks More Effectively?
[…]
$\big\|\mathcal{T}(\tilde{P})\big\|_{\infty}\;\leq\;e^{-\beta}\,\big\|\mathcal{T}(P)\big\|_{\infty}.$ (6)
In parallel, a Jacobian-based estimate provides an additive Lipschitz bound in $\ell_{2}$ for small $\beta$ (App. F.6). Since the cross-domain map in MILD composes $LT$ attention blocks with per-block biases $\{\beta_{\ell,t}\}$, submultiplicativity gives
$\big\|\mathcal{T}_{\mathrm{MILD}}\big\|_{\infty}\;\leq\;\exp\!\Big(-\sum_{\ell=1}^{L}\sum_{t=1}^{T}\beta_{\ell,t}\Big)\,\big\|\mathcal{T}(P_{\mathrm{base}})\big\|_{\infty}\;=\;e^{-\bar{\beta}LT}\,\big\|\mathcal{T}(P_{\mathrm{base}})\big\|_{\infty},$ (7)
where $\bar{\beta}=\tfrac{1}{LT}\sum_{\ell,t}\beta_{\ell,t}$. Eq. 6 and Eq. 7 explain the plots: the erased-instance peak mass admits an upper bound decaying approximately as $e^{-\bar{\beta}LT}$, i.e., linear in depth and steps on a log scale.
5.2 How Does MILD Enable Flexible Scene Recomposition?
Beyond achieving high-quality foreground removal and background restoration, MILD’s layered diffusion design introduces a key advantage: the ability to flexibly recompose scenes using instance-specific foreground layers and a clean background. As shown in Figure 8, MILD generates disentangled outputs for each target, enabling selective combination to construct customized scenes. This allows users to separate specific individuals (e.g., only IP11), remove arbitrary targets (e.g., w/o IP33), or reconstruct any subset of the original scene. Such instance-aware recomposition offers fine-grained control over content manipulation, making MILD well-suited for interactive editing and scene simulation tasks. More details are shown in Appendix G.6.
6 Conclusion
We present Multi-Instance Layered Diffusion, a diffusion-based strategy reformulating inpainting as a layered, decoupled diffusion process for precise human erasing, supported by our curated, high-quality MILD dataset. On top of our theoretical analysis based on CAG, MILD leverages multi-layer structure along with HMG and SMA to provide instance-aware structure understanding and brings spatially-conditioned attention control, which effectively reducing artifacts and improving reconstruction fidelity. Extensive experiments show that MILD outperforms existing inpainting baselines, achieving better semantic consistency, fewer artifacts, and more coherent scene reconstructions. MILD provides a promising direction for scaling precise diffusion-based inpainting to complex multi-object and open-domain removal tasks.
7 Ethics Statement
When gathering our dataset, we ensure that raters are compensated. We also run safety filters over the generated images before giving them to the raters. This work is a step towards better evaluation of text-to-image models which are known to hallucinate. It gives tools to others developers and practitioners to properly understand and evaluate T2I models in the future.
8 Reproducibility Statement
Our project page, along with our code and dataset is available at: project page. We give extensive details of our experimental setup in Section 4.1 and Appendix G.1. For human annotation, we visualise the templates used and give extensive detail on how these raw ratings are aggregated in  G.5. For our dataset , the details are stated in the Appendix A, and the sample data pairs are submitted as supplementary materials. For metrics, we give full details of the baselines in Section 4 and the evaluation pipeline will be released along with code upon acceptance. For model design, we provide the core algorithm in Appendix D, and the code will be released upon acceptance. For theoretical results, clear explanations of any assumptions and a complete proof of the claims are included in the Appendix E and Appendix F.
Appendix A Dataset and Code Availability
Dataset Construction.
Our dataset consists of high-quality images sourced from two main origins: (1) a carefully curated subset of the OpenImages V5 dataset (openimages2020), and (2) publicly available web images, collected under fair use for academic research purposes. All images are manually verified to ensure diversity in scenes, human poses, occlusion patterns, and background complexity.
Figure 9: Sample data collected from MILD dataset.
Each image is annotated with pixel-level, instance-specific human masks. In addition, we provide human-selected ground-truth completions to enable supervised evaluation of object removal performance. Figure 9 presents representative examples from the dataset. The full dataset will be publicly released on our project page upon acceptance, and the sample dataset is provided in our supplementary materials.
Code and Model Availability.
To facilitate future research and ensure reproducibility, we have created an anonymous project page containing representative examples and the paper abstract. Our core codebase and pre-trained models will be updated upon acceptance.
Appendix B The Use of Large Language Models (LLMs)
Large language models (LLMs) were used for grammar and formatting checks, as well as for translation assistance.
Appendix C More Related Works
Image Inpainting
Image inpainting aims to restore missing or occluded regions in a visually coherent and semantically consistent manner in various scenarios (sargsyan2023migan; zhou2020oracle; wang2024gscream; yoon2024corempi). Traditional methods relied on patch propagation (hays2007scene; ding2019nonlocal), transferring low-level textures from visible regions but failing in semantically complex scenes. With the rise of deep learning, the encoder-decoder (propainter2023; cao2023zitspp; altinel2018energy; li2020recurrent; yi2020contextual) and GAN-based architectures (iizuka2017globally; yu2018generative; zeng2019learning) have significantly improved structural plausibility and perceptual quality, further enhanced by edge guidance (nazeri2019edgeconnect) and attention mechanisms (yu2018generative). Diffusion-based methods (lugmayr2022repaint; saharia2022palette) have recently achieved strong generation performance by operating in latent spaces guided by powerful pretrained priors, but they often lack precise spatial control or instance-level disentanglement. Methods like SDXL Inpainting (podell2023sdxl) introduce mask-based conditioning for edit localization, but treating all masked regions as single entity prevents instance-specific control. Meanwhile, recent efforts incorporate large language models (LLMs) to enable instruction-based editing (brooks2023instructpix2pix; huang2024smartedit; li2024zone; sheynin2024emuedit; li2024gligen), yet they struggle to localize individual instances in complex scenes.
Layered Generation and Controllability.
Recent works on image generation have explored layered modeling and layout-based generation to enhance object awareness and controllability. Layered modeling works (jia2024designedit; lu2020layered) leverage multi-level structures through latent decomposition and scene-layer separation to enable semantically consistent editing. However, they often operate within a unified generation pathway and lack explicit instance disentanglement or spatial conditioning, making them less suitable for multi-object removal. Prompt- and layout-controllable generation works (liu2022composable; zheng2023layoutdiffusion; zhangli2024layout) aim to synthesize scenes from object-level prompts or layout priors. Although these methods offer flexible conditioning and structural guidance, they generally operate at a holistic level without explicit separation of foreground instances, limiting their capacity to handle occlusions or support fine-grained instance control. In contrast, our approach introduces a mask-aware, multi-branch generation framework that explicitly disentangles each foreground instance and the background into independent pathways. This layered formulation enables precise object removal, instance-level recomposition, and enhanced semantic fidelity through joint supervision.
Appendix D More Algorithmic Details
To ensure reproducibility, we provide additional details of the training pipeline, architectural design, and spatial supervision strategies used in LayeredUNet. The model is optimized in stages (Algorithm 2), where a training controller gradually activates different generation branches. The training detail is in Algorithm 1. Each image is encoded into the latent space and perturbed with spatial offsets derived from the union of all masks. To stabilize learning, the background branch is initially protected by freezing its gradients, while foreground branches are progressively optimized with instance-specific conditioning.
@@ Appendix D More Algorithmic Details
14:   $\mathcal{L}_{k}\leftarrow\|M_{k}\odot(\epsilon-\epsilon_{k})\|_{2}^{2}$ {Foreground branch}